### 08 Jan 2024 v1

# Is Limited Information Enough? An Approximate Multi-agent Coverage Control in Non-Convex Discrete Environments

Abstract.
Conventional distributed approaches to coverage control may suffer from lack of convergence and poor performance, due to the fact that agents have limited information, especially in non-convex discrete environments. To address this issue, we extend the approach of Marden (2016) which demonstrates how a limited degree of inter-agent communication can be exploited to overcome such pitfalls in one-dimensional discrete environments. The focus of this paper is on extending such results to general dimensional settings. We show that the extension is convergent and keeps the approximation ratio of 2, meaning that any stable solution is guaranteed to have a performance within 50% of the optimal one. We also show that the computational complexity and communication complexity are both polynomial in the size of the problem. The experimental results exhibit that our algorithm outperforms several state-of-the-art algorithms, and also that the runtime is scalable as per theory.
Key words and phrases:
Multi-agent systems; Coverage control; Communication
1. Introduction
Coverage control is a fundamental problem in the field of multiagent systems (MAS). The objective of a coverage control problem is to deploy homogeneous agents to maximize a given objective function, which basically captures how distant the group of agents as a whole is from a pre-defined set of Points of Interest (PoI). Coverage control has a wide range of applications, such as tracking, mobile sensing networks or formation control of autonomous mobile robots Cortes et al. (2004).
It is known that, even in a centralized context, finding an optimal solution for the coverage problem is an NP-hard problem Megiddo and Supowit (1984). Hence, most studies focus on approximate approaches. In distributed settings, game-theoretical control approaches seek to design agents that will be incentivized to behave autonomously in a way that is well-aligned with the designer’s objective. This strategy has proven to be successful in a number of applications (see, e.g. Douchan et al. (2019)). The situation is even more difficult in practice since agents often have restricted sensing and communication capabilities. Consequently, agents must make decisions based on local information about their environment and the other agents. Unfortunately, algorithms based on local information may suffer from lack of convergence and degraded performance due to miscoordination between the agents. As for the convergence issue, it is known that a move of an agent can affect the cost of agents outside of the neighborhood, and thus, the decrease of the global cost cannot be guaranteed locally, especially in the case of discrete non-convex environment Yun and Rus (2014). The degradation of performance can be explained with a worst-case scenario in which only a single agent can perceive a large number of valuable locations within an environment, while a number of other agents cannot perceive these locations.
Recently, Marden Marden (2016) made more precise the connection between the degree of locality related to the available information and the achievable efficiency guarantees. He showed that the achievable approximation ratio depends on the individual amount of information available to the agents. Consequently, distributed algorithms are inevitably subjected to poor worst-case guarantees because of the locality of the information used to make decisions. If all agents have full global information as in the case of centralized control, there exist decentralized algorithms that give a 2 approximation ratio. Conversely, under limited information (e.g. Voronoi partitions), no such algorithm provides such an approximation ratio. Rather, the best decentralized algorithm that adheres to these informational dependencies achieves, at most, an $n$ approximation factor, where $n$ is the number of agents. Then, the focus in MAS settings is on how to design agents that, through sharing a limited amount of information with neighborhood communications, achieve an approximation ratio close to 2.
Indeed, different settings exist that vary according to the information available:
  1. (1)
    agents may not communicate any information and be only guided by their local perception;
  2. (2)
    agents may communicate candidate individual moves to their neighbors only;
  3. (3)
    agents may communicate candidate moves (possibly coordinated with other agents) to agents (possibly beyond their neighbors).
Existing game-theoretical approaches can be classified into these types: classical Voronoi-based best-response approaches fall into either (1) or (2), depending on the assumption. Two recent approaches in type (3) explore different directions. Sadeghi et al. Sadeghi et al. (2020) proposed a distributed algorithm for non-convex discrete settings in which agents have the possibility to coordinate a move with a single other agent (meaning that an agent moves and assumes another agent takes her place simultaneously), possibly beyond their neighbors, when individual moves are not sufficient. However, the algorithm sacrifices the approximation ratio for convergence. Marden Marden (2016) on the other hand allows coordinated moves with several agents, but only under the restriction that these agents are within the neighborhood. While this approach achieves a convergent algorithm with an approximation ratio of 2, by sharing only the information of the minimum utility among agents, it is limited in that the only investigated case is the one-dimensional (line) environment. This severely limits real-world applications. Other studies try to achieve global optimality by developing approaches akin to simulated annealing. For example, Arslan et al. (2007) attains global optimum with Spatial Adaptive Play (SAP, a.k.a Boltzmann exploration) and uses random search to escape from the local optimum. However, this approach suffers from a slow convergence rate when the search space is large. There is also no discussion about the relationship between information and efficiency. Note that we focus on the coverage problem, and multiagent path-planning issues Galceran and Carreras (2013); Stern et al. (2019); Okumura and Défago (2022) are out of the scope of the paper.
In this paper, we extend the algorithm of Marden (2016) to any-dimensional, non-convex discrete space and compare this approach with the aforementioned alternative variants of game-theoretical control. Our approach can also be considered as a generalization of Durham et al. (2011), while they are limited to pairwise neighborhood without a guarantee of approximation ratio. The remainder of this paper is as follows. Section 2 introduces the model and existing approaches. In Section 3 we detail the algorithm and prove that it guarantees convergence to a neighborhood optimum solution with an approximation ratio of 2 without any restriction on the dimensionality of the environment, i.e., the same guarantee as in the 1D case. Additionally, we propose a polynomial-time extension of the algorithm and discuss the computation and communication complexity. Experiments reported in Section 4 indeed show that our algorithm outperforms existing ones, along with adhering to the theoretical approximation ratio. Furthermore, the runtime results confirm the scalability of the proposed algorithm.
2. Model
2.1. Coverage Problems
We start with a set of agents $\mathcal{N}=\{1,\ldots,n\}$ and a set of resources $C=\{c_{1},\ldots,c_{m}\}$. In a coverage problem, resources are locations (or points) in a connected metric space. We assume that this is discrete finite space that is modeled as a connected graph $(C,\mathcal{E})$, where $\mathcal{E}$ is the set of edges that connect two adjacent points. Though our approach can be extended to continuous settings, we omit the detail due to the space limit. We denote the distance between two points $a,b\in C$ as $|a-b|$ that is the length of the shortest path.
An allocation $x$ maps each agent $i$ in $\mathcal{N}$ to a resource (i.e., a point) in $C$. An allocation is thus defined as a vector of resources $x=\langle x_{i}\in C|i\in\mathcal{N}\rangle$ where $x_{i}$ is the resource assigned to agent $i$. Note that each agent must be allocated one and only one resource. An allocation is exclusive i.e., $x_{i}\neq x_{j}$ $\forall i,j\in\mathcal{N}$ such that $i\neq j$. We denote the set of all possible allocations as $\mathcal{X}=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\times}\mathcal{X}_{i}$ where $\mathcal{X}_{i}$ is the set of possible positions for agent $i$.
Let $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ denote a non-increasing function, $v_{c}\in\mathbb{R}_{+}$ be the weight of point $c\in C$, and $\mathcal{C}\subseteq C$ be a partial space. Then, the objective function for $\mathcal{C}$ is defined as follows:
$\left.\begin{array}[]{l}G(x;\mathcal{C})=\underset{\begin{subarray}{c}c\in\mathcal{C}\end{subarray}}{\sum}\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}\leavevmode\nobreak\ v_{c}\leavevmode\nobreak\ g(|x_{i}-c|).\end{array}\right.$ (1)
For simplicity, we denote $G(x)=G(x,C)$. The goal of the coverage problem is then to find an optimal allocation $x^{*}\in\mathcal{X}$ such that:
$\left.\begin{array}[]{l}x^{*}\in\underset{\begin{subarray}{c}x\in\mathcal{X}\end{subarray}}{\mathop{\rm argmax\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ G(x).\end{array}\right.$ (2)
[…]
Figure 1. Coverage example with 6 agents: Circles in grey are agents. A number in a white circle shows the Manhattan distance to the closest agent.
Example 0.
Let us then consider the coverage problem depicted in Figure 1. For the sake of exposure, let us assume that $g(d)=1/(1+d)$, and ||$\cdot$|| ⋅ | is Manhattan distance. Agents are identified by letters $a,b,\dots f$. The environment is a grid world, where circled locations are valued $v_{c}=1$, while the others are valued $v_{c}=0$. The locations covered by an agent are represented in grey and we indicate the name of the agent. For unoccupied locations we indicate on the node the (Manhattan) distance to the closest agent. Thus a covered location gives utility of 1, while a location at a distance of 1 from the closest agent gives utility 1/2, and so on. The depicted allocation $x$ has value $G(x)=1+1+\frac{1}{2}+\dots=16.4$. It is clearly sub-optimal. The reader can check that the allocation $x^{\prime}$ where agent $e$ has moved one location up would induce $G(x^{\prime})\simeq 16.8$.
It is well-known that solving optimally this problem is NP-hard Megiddo and Supowit (1984). Nevertheless, thanks to the submodular nature of the objective function $G$, the centralized greedy algorithm which allocates agents one by one (starting from the empty environment) guarantees $1-1/e\approx$ 63% of optimal Calinescu et al. (2011). We will use this algorithm as a baseline for comparison.
2.2. Game-Theoretic Control: Generalities
Since we focus on distributed algorithms, each agent has to make a choice about her position based on the partial information she has about the other agents. We thus adopt a game theoretic approach where each agent computes her best response based on her individual information Nisan et al. (2007); Shoham and Leyton-Brown (2008).
In the following, $-i$ will denote the set of agents excluding $i$: $-i=\mathcal{N}\setminus\{i\}$. A partial allocation for a subset of agents $S\subseteq\mathcal{N}$ will be denoted as $x_{S}=\{x_{i}\}_{i\in S}$.
This paper will focus on providing each agent $i$ with a utility function of the form $u_{i}:\mathcal{X}\to\mathbb{R}$ that will ultimately guide their individual behavior. In the rest of the paper, we formulate the set of choices of agent $i$ as $\mathcal{X}_{i}(x_{-i})$ assuming that it depends on a given allocation of the other agents $x_{-i}$. For simplicity, we denote $\mathcal{X}_{i}(x)=\mathcal{X}_{i}(x_{-i})$. We also denote $\mathcal{X}(x)=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\times}\mathcal{X}_{i}(x)$.
When each agent $i\in\mathcal{N}$ selects the position $x_{i}\in\mathcal{X}_{i}(x)$ that maximizes her utility given the other agent’s positions $x_{-i}$, the resulting allocation $x=\langle x_{1}\cdots x_{n}\rangle$ can be a (pure) Nash equilibrium such that:
$\left.\begin{array}[]{l}\forall i\in\mathcal{N}\textrm{\leavevmode\nobreak\ }u_{i}((x_{i},x_{-i}))\geq u_{i}((x_{i}^{\prime},x_{-i})),\forall x_{i}^{\prime}\in\mathcal{X}_{i}(x),x_{i}\neq x_{i}^{\prime}.\end{array}\right.$ (3)
In general, the efficiency of a Nash equilibrium $G(x)$ can be smaller than the optimal value $G(x^{*})$. One of the reasons of this suboptimality is the miscoordination among self-interested agents, as agents are required to make independent decisions in response to available information. Other sources of suboptimality comes from the structure of the agents’ utility functions and available choice sets. The (worst-case, among all instances) ratio between the worst Nash equilibria and the social optimum is known as the price of anarchy (PoA) Koutsoupias and Papadimitriou (1999). Here, the choice set $\mathcal{X}_{i}$ could encode physical constraints on choices, i.e. an agent can only physically choose a given subset of choices given the behavior of the collective $x$. Alternatively, the choice set could encode information availability of the agent, e.g. it is the set of choices for which the agent can evaluate its utility. Regardless of the interpretation, it is important to highlight that the structure of the choice sets can significantly alter the structure and efficiency of the resulting equilibria which we discuss in the ensuing section.
2.3. Local Information
Several approaches in game theory seek to exploit the fact that agents typically have limited sensing power and only a local view of the situation. Marden Marden (2016) analyzed how miscoordination among agents with limited information leads to inefficient Nash equilibria. To this end, he introduced the concept of information set which is the set of choices each agent can perceive and compute the resulting utilities. In this paper, $\mathcal{X}_{i}(x)$ corresponds to the information set that is the set of locations agent $i$ can perceive based on spatial proximity. With this notion, an allocation is a Nash equilibrium (Equation (3)) if agents are at least as happy as any choice for which they can evaluate their utility. Observe that the information set $\mathcal{X}_{i}(x)$ is a state-dependent notion: the local information available may vary depending on the current allocation $x$.
To model to what extent the information is localized, Marden defined the following redundancy index associated to the agents’ local information sets $\{\mathcal{X}_{i}\}_{i\in\mathcal{N}}$:
$\left.\begin{array}[]{l}f=\underset{\begin{subarray}{c}x\in\mathcal{X}\end{subarray}}{\min}\underset{\begin{subarray}{c}y\in C\end{subarray}}{\textrm{\leavevmode\nobreak\ }\min}\textrm{\leavevmode\nobreak\ }|\{i\in\mathcal{N}:y\in\mathcal{X}_{i}(x)\}|.\end{array}\right.$ (4)
Intuitively, $f$ represents the minimum number of agents that perceive the same resource available for their choice. In particular, we note that:
  • $f>0$ guarantees that all the locations are always a possible choice for some agent of the system;
  • If there is an allocation $x$ where a resource is a possible choice for only one agent and all other resources are possible choices for at least one agent, then $f=1$.
  • $f=n$ is the extreme case of full information access where all resources are possible choices for all agents.
2.4. Linking Information and Inefficiency
Intuitively, local information can cause distributed systems based on game-theoretic control to get stuck in an inefficient allocation. Indeed, some agents who may obtain higher utilities for a resource may not have access to this information. The redundancy index hence gives insights about the amount of information available to the agents and guarantees on the worst-case ratio.
Marden Marden (2016) investigated this interplay, in the broader context of resource allocation games. Assuming that the global objective function $G$ is monotone submodular i.e., satisfies the following conditions:
$\left.\begin{array}[]{l}G(x_{T})\geq G(x_{S}),\forall S\subseteq T\subseteq N.\\ G(x_{S})-G(x_{S\setminus\{i\}})\geq G(x_{T})-G(x_{T\setminus\{i\}}),\forall S\subseteq T\subseteq\mathcal{N},\forall i\in S.\\ \end{array}\right.$ (5)
and satisfies two further properties. First, the utility of each agent is greater than her marginal contribution:
$\left.\begin{array}[]{l}u_{i}(x)\geq G(x)-G(x_{-i}).\end{array}\right.$ (6)
Second, social welfare is less than the global objective:
$\left.\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x)\leq G(x).\end{array}\right.$ (7)
In the later section, we will see that the global objective function and the utility function of the conventional coverage control both satisfy these assumptions.
The following theorem shows how the value of $f$ impacts the efficiency of Nash equilibrium allocation:
Theorem 2 (from Marden (2016)).
If $G$ is a monotone submodular set function and $u_{i}$ satisfies (6) and (7), the worst case efficiency of Nash equilibrium $x$ is lower bounded by
$\left.\begin{array}[]{l}G(x)\geq\frac{f}{n+f}G(x^{*}).\end{array}\right.$ (8)
Furthermore, there exists a case such that:
$\left.\begin{array}[]{l}G(x)=\frac{f}{n}G(x^{*}).\end{array}\right.$ (9)
Thus, we know that the approximation ratio of a distributed allocation algorithm can be 2 in the case of full information ($f=n$), because $G(x)\geq\frac{1}{2}G(x^{*})$ in this case (from Equation (8)). Also, it is impossible to guarantee an approximation ratio better than $n$ in case of $f=1$ (from Equation (9)). We will use this result in the later section to analyze the efficiency of coverage control algorithms.
2.5. Voronoi-Based Control
One of the most standard approaches of coverage control is the algorithms based on Voronoi partitioning Gao et al. (2008). A Voronoi partition divides the space into local regions for each agent. Formally, for a given allocation $x$, a Voronoi partition $\mathcal{V}_{i}(x;\mathcal{C})$ of space $\mathcal{C}$ is defined as follows:
$\left.\begin{array}[]{l}\mathcal{V}_{i}(x;\mathcal{C})=\{c\in\mathcal{C}|i=\underset{\begin{subarray}{c}j\in\mathcal{N}\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x_{j}-c|\}.\par\end{array}\right.$ (10)
When ties occur (i.e., when several agents are at the same minimal distance from some location), agents are prioritized lexicographically.
We assume that the locations an agent can perceive are limited to those within their Voronoi region, that is $\mathcal{X}_{i}(x)=\mathcal{V}_{i}(x;C)$. We also denote the neighborhood of $i$ as the agents connected in the dual Delaunay graph of the Voronoi partition of $i$, i.e., the neighborhood of $\mathcal{N}_{i}\subseteq\mathcal{N}$ are the agents $j\in\mathcal{N}$ whose Voronoi region is connected to the Voronoi region of $i$.
The utility function of agent $i$ is then defined as follows:
$\left.\begin{array}[]{l}u_{i}(x)=\underset{\begin{subarray}{c}c\in\mathcal{X}_{i}(x)\end{subarray}}{\sum}v_{c}\leavevmode\nobreak\ g(|x_{i}-c|).\end{array}\right.$ (11)
Note that these utility functions satisfy the assumptions of Equation (6) and Equation (7). Voronoi partition plays an important role in distributed coverage control because agents can compute the best responses improving the objective function locally, with limited communication. The process is then just a sequence of best response updates (in the sense defined above) of the different agents to compute their next locations in their local Voronoi region. Once agents are assigned these new locations, Voronoi regions are updated and the process iterates, until convergence. However, in the non-convex discrete setting the move of an agent within its partition can affect not only neighbors (as in the continuous setting), but the whole set of agents in the worst case Yun and Rus (2014). (We will see an example later in Figure 5). In theory, the guarantee of convergence requires avoiding moves that could impact beyond an agent’s neighbors Yun and Rus (2014); Sadeghi et al. (2020).
Observe that the Voronoi partition induces a redundancy index of $f=1$ because all the points in Voronoi regions $\mathcal{X}_{i}$ are available only for agent $i$. Then, the efficiency $G(x)$ can be $1/n$ of the optimum value in the worst case.
[…]
$a$$b$$c$$d$$e$$f$
Figure 2. The environment of Example 1 with Voronoi partitions and the corresponding neighborhood graph
Example 0.
(Ex. 1, cont.). Figure 2 indicates the Voronoi region of each agent by coloring the locations in the same color. (Recall that ties are broken lexicographically). Figure 2 gives the corresponding neighboring graph: agent $e$ has 3 neighbors $\{c,d,f\}$. The utilities of the agents are as follows: $u_{a}=1$, $u_{b}=1.5$, $u_{c}\simeq 3.2$ $u_{d}=4.2$ $u_{e}\simeq 5$ and $u_{f}=1.5$
3. A Neighborhood Optimal Algorithm
To address the inefficiency issue of coverage control through inter-agent communication, we extend the solution for 1-dimensional space Marden (2016) to general settings by making agents share additional information. We first present our main result on convergence and approximation ratio and then discuss the details of the algorithm.
3.1. High Level Description of the Algoritm
The idea of the algorithm is to incrementally compute an allocation $x$ based on a partition $\mathcal{P}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{N})$ of the space, which is not necessarily a Voronoi partition. Note that the neighborhood $\mathcal{N}_{i}$ is also defined over $\mathcal{P}$, instead of Voronoi partition. The algorithm is anytime and updates a solution $(x,\mathcal{P})$ for each iteration.
As in the case of Voronoi-based control, agents’ utilities $u_{i}(x,\mathcal{P})$ are localized and depend on the current partitioning of the space, i.e., we have $\mathcal{X}_{i}(x)=\mathcal{P}_{i}(x;C)$. The minimum utility enjoyed by any agent of the system is $u_{\min}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\min}\leavevmode\nobreak\ u_{i}(x,\mathcal{P})$, and the agent with this minimum utility is denoted as $i_{\min}(x,\mathcal{P})$. When there are multiple candidates for $i_{\min}$ it can be resolved using a tie-breaking rule. We first discuss the case of single $i_{\min}$. The case of multiple candidates of $i_{\min}$ is discussed later.
We follow Marden Marden (2016) and define the maximum gain in $G$ when adding $k$ new agents into space $\mathcal{C}$, as:
$\left.\begin{array}[]{l}M_{k}(x,\mathcal{C})=\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{C}\end{subarray}}{\max}G(y_{1},\ldots,y_{k},x;\mathcal{C})-G(x;\mathcal{C}),\end{array}\right.$ (12)
where this best allocation of the $k$ new agents is denoted as follows:
$\left.\begin{array}[]{l}B_{k}(x,\mathcal{C})=\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{C}\end{subarray}}{\mathop{\rm argmax\leavevmode\nobreak\ }\limits}G(y_{1},\ldots,y_{k},x;\mathcal{C})-G(x;\mathcal{C}).\end{array}\right.$ (13)
Note that determining the locations $y_{1},\dots y_{k}$ maximizing Equation (13) is actually an NP-hard problem, since it consists in general in solving a multiagent coverage problem.
Given this, we can find the agent $i_{\max}^{+}$ which would contribute to $G(x)$ the most from having an additional agent in her region:
$\left.\begin{array}[]{l}i_{\max}^{+}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\mathop{\rm argmax\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ M_{1}(x_{i};\mathcal{P}_{i}).\end{array}\right.$ (14)
$\left.\begin{array}[]{l}V(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}\leavevmode\nobreak\ M_{1}(x_{i};\mathcal{P}_{i}),\end{array}\right.$ (15)
Example 0.
(Ex. 1, cont.). We have $i_{min}=a$, since $u_{a}=1$. Furthermore, the agent which would contribute to $G(x)$ the most from adding a single agent within her region is agent $e$, as adding an agent (depicted as ’+’) would induce $G$ of $6.5$ in her region, thus $M_{1}(x_{e},\mathcal{P}_{e})\simeq 6.5-5=1.5$ (See Figure 3 for the illustration).
[…]
Figure 3. Agent $e$ would contribute to $G(x)$ the most from the addition of an agent (depicted as ’+’) in her region
Before going into the details, we show the following main result.
Theorem 5 (Convergence with Performance guarantee).
Assuming that agents have access to $u_{\min}$ and $i_{\max}^{+}$ given any configuration, Algorithm 2 converges to a neighborhood optimum and terminates. The approximation ratio of the algorithm is 2.
Note that the approximation ratio of 2 is equal to the lower bound predicted by Equation (8) in case of full information ($f=n$). Now we explain how we construct Algorithm 2 and reach this conclusion. We classify the solution $(x,\mathcal{P})$ into 4 states. First, the solution space is divided into the following two states $Z_{1}$ and $Z_{2}$, by checking if $i_{\max}^{+}$ would gain more from adding an agent in her region:
$Z_{1}=$$\{(x,\mathcal{P})|V(x,\mathcal{P})>u_{\min}(x,\mathcal{P})\},$ (16)
$Z_{2}=$$\{(x,\mathcal{P})|V(x,\mathcal{P})\leq u_{\min}(x,\mathcal{P})\}.$ (17)
In $Z_{2}$, no single agent would contribute enough from accommodating $i_{\min}$. Next the solutions in $Z_{2}$ are further classified depending on whether integrating a further agent in the neighborhood could induce a significant enough marginal gain. We denote $n_{i}=|\mathcal{N}_{i}|$ and $\mathcal{P}_{n_{i}}=\cup_{j\in\mathcal{N}_{i}}\mathcal{P}_{j}$. In the following state $Z_{3}$, there is no such a group:
$Z_{3}=$$\{(x,\mathcal{P})\subseteq Z_{2}|$
$M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P}),$ (18)
$\leavevmode\nobreak\ \forall i\in\mathcal{N}\},$
Intuitively, $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})$ is the efficiency of a neighborhood optimum of $\mathcal{N}_{i}$. Equation (18) means that the gain in the neighborhood optimum by accommodating $i_{\min}$ in $\mathcal{P}_{n_{i}}$ cannot be larger than $u_{\min}$.
Lastly, a solution is classified as state $Z_{4}$ if neighborhood optimality Weiss and Freeman (2001) is achieved for all agents. $Z_{4}$ is the terminal state that is reached from a solution in $Z_{3}$ if it satisfies the following condition:
$Z_{4}=$$\{(x,\mathcal{P})\subseteq Z_{3}|$
$\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})=M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall i\in\mathcal{N}\}.$ (20)
Example 0.
(Ex. 1, cont.) The allocation $x$ depicted is in state $Z_{1}$ since $V(x,\mathcal{P})=1.5>1=u_{a}$, with $i_{min}=a$.
Algorithm 1 Communication tree
1:procedure $D_{i}=$CommTree($i$)
2:     $u_{{\rm lmin}}=\hat{u}_{i}$, $j_{{\rm lmin}}=i$
3:     for $j\in\mathcal{\hat{N}}_{i}$  do
4:         if $\hat{u}_{j}<u_{{\rm lmin}}$ then
5:              $u_{{\rm lmin}}=\hat{u}_{j}$, $j_{{\rm lmin}}=j$
6:     $f_{updated}=False$
7:     if $j_{{\rm lmin}}=i$  then
8:         if $D_{i}<0$ then
9:              $D_{i}=i$, $f_{updated}=True$
10:     else
11:         if $D_{i}\not=j_{{\rm lmin}}$ then
12:              $D_{i}=j_{{\rm lmin}}$
13:              $\hat{u}_{i}=u_{{\rm lmin}}$, $f_{updated}=True$
14:     if $f_{updated}$ then
15:         for $j\in\mathcal{\hat{N}}_{i}$  do
16:              $D_{j}=\textsc{CommTree}(j)$
17:     Return $D_{i}$
[…]
Figure 4. (Top left): the neighborhood graph in Example 1. The numbers represent utilities. (Top right): agent ’c’ points to the agent ’b’ that has minimum utility among neighborhood. It also updates its utility. (Bottom left): then agents ’b’ and ’e’ point to the minimum neighbors. (Bottom right): the communication tree in convergence. In this case, agent ’a’ is $i_{\min}$ and therefore the root of the tree.
Algorithm 2 Neighborhood optimum algorithm
1:procedure $x=$NeighborOpt($x,\mathcal{C}$)
2:     $\mathcal{P}_{i}=\mathcal{V}_{i}(x;\mathcal{C}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$
3:     while $(x,\mathcal{P})\not\in Z_{4}$  do
4:         Communicate $u_{\min}(x,\mathcal{P})$, $x_{i_{\min}}(x,\mathcal{P})$ and $i_{\max}^{+}(x,\mathcal{P})$
5:         Pick $i\in\mathcal{N}$
6:         $\mathcal{N}^{2}_{i}=\textsc{Neighbor}(\mathcal{P},i)$
7:         while $(x,\mathcal{P})\in Z_{1}$ and $i_{\max}^{+}\not\in\mathcal{N}^{2}_{i}$  do
8:              pick $i\in\mathcal{N}$
9:         if $i_{\min}(x,\mathcal{P})\in\mathcal{N}^{2}_{i}$ or
10:            $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P})$  then
11:               $x\leftarrow B_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})$ (Step a)
12:              $\mathcal{P}_{j}\leftarrow\mathcal{V}_{j}(x^{\prime},\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall j\in\mathcal{N}^{2}_{i}$
13:         else
14:              Consider a virtual agent $l\not\in\mathcal{N}$ (Step b)
15:              $x^{\prime}\leftarrow B_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})$
16:              $\mathcal{P}^{\prime}_{k}\leftarrow\mathcal{V}_{k}(x^{\prime},\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall k\in\mathcal{N}^{\prime}_{i}=\{1,\ldots,n_{i}+1\}$
17:              $\tilde{i}_{\min}=\underset{\begin{subarray}{c}j\in\mathcal{N}_{\mathcal{N}^{2}_{i}}\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x_{j}-x_{i_{\min}}|$
18:              $k_{l}=\underset{\begin{subarray}{c}k\in\mathcal{N}^{\prime}_{i}|{\rm adj}(\mathcal{P}^{\prime}_{k},\mathcal{P}_{\tilde{i}_{\min}})\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x^{\prime}_{k}-x_{\tilde{i}_{\min}}|,\leavevmode\nobreak\ x_{l}\leftarrow x^{\prime}_{k_{l}}$
19:              $x\leftarrow x^{\prime}\setminus\{x_{l}\}$
20:              $i_{+}\leftarrow\underset{\begin{subarray}{c}j\in\mathcal{N}^{2}_{i}|{\rm adj}(\mathcal{P}^{\prime}_{j},\mathcal{P}^{\prime}_{l})\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ |x_{j}-x_{l}|,\leavevmode\nobreak\ \mathcal{P}_{i_{+}}\leftarrow\mathcal{P}_{i_{+}}\cup\mathcal{P}_{l}$
21:     Return $(x,\mathcal{P})$
Observe that these computations need to broadcast the global information of $u_{\min}(x,\mathcal{P})$, $x_{i_{\min}}(x,\mathcal{P})$ and $i_{\max}^{+}(x,\mathcal{P})$. For this purpose, we propose an algorithm based on message exchanges among neighbors. Agent $i$’s neighborhood $\mathcal{N}_{i}$ is defined as the set of agents whose partitions are adjacent to $\mathcal{P}_{i}$ as follows:
$\mathcal{N}_{i}=\{j\in\mathcal{N}\setminus\{i\}|{\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})\},$ (21)
where ${\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})=(\exists(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j})$ returns true if $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ are adjacent.
Algorithm 1 shows how to share global information, for example, $u_{\min}$, by constructing a communication tree in a distributed manner. The algorithm can be triggered by any agent $i$. The goal of the algorithm is to find agent $D_{i}$ that is the only parent of $i$ in the communication tree. Prior to executing the algorithm, each agent initializes its memorized utility $\hat{u}_{i}$ with its own utility $u_{i}$. Also, we set $\mathcal{\hat{N}}_{i}=\mathcal{N}_{i}$. Then firstly, agent $i$ communicates with its neighbors $\mathcal{\hat{N}}_{i}$ to find the agent $j_{{\rm lmin}}$ that has the locally minimum utility among them (Lines 2-5). If $i$ itself is $j_{{\rm lmin}}$, $i$ points to itself only when it has not pointed any other agents (Lines 7-9). If $j_{{\rm lmin}}$ is not either $i$ nor the current $D_{i}$, $i$ points to $j_{{\rm lmin}}$ (Lines 10-12). At the same time, $i$ also updates its utility in memory to serve as a proxy to the root ($i_{\min}$) in further communications (Line 13). Lastly, the algorithm of the neighborhood is triggered recursively until the convergence (Line 16). In the convergence, a tree rooted at $i_{\min}$ is formed through the chain of pointers $D_{i}$.
Figure 4 shows how the communication tree is constructed in the case of Example 1. In convergence, all the agents in the tree have a consensus on $i_{\min}$ (and $u_{\min}$), which is the root of the tree. The communication tree plays an important role not only in sharing global information but also in algorithm scalability. As we mentioned, the computation of Equation (13) is NP-hard and is not scalable in the size of the neighborhood $n_{i}$. To address this issue, we define a fixed size neighborhood, $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$, where each agent $i$ has only its parent $D_{i}$ as a single neighbor and therefore $n_{i}=2$. We also update the definition of adjacency function to maintain consistency with $\mathcal{N}^{2}_{i}$ as ${\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})=(\exists(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j},j\in\mathcal{N}^{2}_{i})$. The distance function ||$\cdot$|| ⋅ | also needs to be redefined only on the edges in the communication tree, $\mathcal{E}\setminus\{(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j},\neg{\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})\}$.
Given these notations, we propose a distributed algorithm returning an equilibrium solution $(x,\mathcal{P})$. The algorithm is summarized in Algorithm 2. At each iteration, an agent is activated. The agent communicates and coordinates with her neighborhood to approach a neighborhood optimum. Firstly, all agents initialize $\mathcal{P}_{i}$ with a Voronoi partition based on geodesic distances Haumann et al. (2011) (Line 2). As far as the state is not $Z_{4}$, each agent iterates the algorithm and shares the necessary information, $u_{\min}$, $x_{i_{\min}}$ and $i_{\max}^{+}$ based on a consensus algorithm via neighborhood communication, as in Algorithm 1 (Lines 3-4).
Then the algorithm picks an agent $i$ (Line 5) in a distributed fashion. This is, for instance, done by sharing ’done/undone’ status and agent IDs, to pick up the ’undone’ agent with the smallest ID. Alternatively, agents could probabilistically move (probability $\alpha$) or not (probability $1-\alpha$). Agent $i$ computes its neighborhood $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$ (Line 6). If the state is $Z_{1}$, it continues to pick another agent until $i_{\max}^{+}$ is in $\mathcal{N}^{2}_{i}$ to prioritize the process of $i_{\max}^{+}$ (Line 7-8). The activated agent $i$ then checks whether the neighborhood $\mathcal{P}_{i}$ could accommodate another agent (Line 10).
  • Step a: If $\mathcal{P}_{n_{i}}$ cannot accommodate another agent, agent $i$ computes a neighborhood optimum for $\mathcal{N}^{2}_{i}$ and implements the allocation (Line 11).
  • Step b: If $\mathcal{P}_{n_{i}}$ can accomodate another agent, it computes new locations $x^{\prime}$ for agents $\mathcal{N}^{2}_{i}$ with an additional agent $l$ (Line 15). Then, the algorithm makes a new partition for $n_{i}+1=3$ agents by splitting $\mathcal{P}_{n_{i}}$ based on the new locations $x^{\prime}$ (Line 16). To allocate the new partition, the algorithm finds agent $\tilde{i}_{\min}\not\in\mathcal{N}^{2}_{i}$ that is the closest to $i_{\min}$ in $\mathcal{N}_{\mathcal{N}^{2}_{i}}$, that is the neighbors of $\mathcal{N}^{2}_{i}$ (Line 17). Then the partition closest to $\tilde{i}_{\min}$ and adjacent to $\mathcal{P}_{\tilde{i}_{\min}}$ is allocated to $l$ first (Line 18). The other partitions are allocated to $\mathcal{N}^{2}_{i}$ by, for example, the optimal transport algorithm (Line 19). Lastly, $\mathcal{P}_{l}$ is merged to the partition of the agent that is the closest and adjacent to $l$ (Line 20).
[…]
Figure 5. A pathological non-convex discrete example. Different from grid spaces, nodes are connected by edges. Left: Voronoi partition. Middle: A move of agent ’a’ changes partitions outside of the neighborhood. Right: Changes in $\mathcal{P}$ are confined inside the neighborhood.
In both cases, agents allocate the partition to avoid collisions after they redevide the neighborhood. Note that the algorithm updates the partition $\mathcal{P}_{j}$ of only neighbors $j\in\mathcal{N}^{2}_{i}$ and does not affect the other agents outside of $\mathcal{N}^{2}_{i}$ (Figure 5). This guarantees the convergence of the algorithm even in non-convex discrete settings as in Theorem 5.
The complete proof of Theorem 5 is in the supplementary material iwa (2024) due to the space limit. Briefly, it proves that the following potential function always increases for each iteration of the algorithm. Since the solution space is finite or compact, then the algorithm terminates.
$\left.\begin{array}[]{l}\phi(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{.}\end{array}\right.$ (22)
Also, the proof sketch of the approximation ratio is as follows. Because of the submodularity of $G$ and Equation (18), adding agents in optimal allocation $x^{*}$ to the same number of agents allocated by the algorithm does not make $G$ double. Formally, $G(x^{*})\leq G(x,x^{*})\leq 2G(x)$.
As shown in Theorem 5, the algorithm achieves the approximation ratio of 2, by communicating the minimum degree of information, $u_{\min}$, $i_{\max}^{+}$ and $x_{i_{\min}}$. The agent with the maximal gain from an additional agent in his region $i_{\max}^{+}$ is used to focus on the area to be reallocated in Line 7 of Algorithm 2. The minimum utility $u_{\min}$ is the key to proving the approximation ratio based on Equation (18). The location $x_{i_{\min}}$ is required to extend the algorithm for 1-dimensional setting in Marden (2016) to more general settings. Note that Algorithm 2 is not the only one that achieves the approximation ratio of 2. The required information also depends on the algorithm and $u_{\min}$, $i_{\max}^{+}$ and $x_{i_{\min}}$ are not the necessary conditions of the approximation ratio.
The algorithm works even better in a special case. Let $\mathcal{C}_{+}\subseteq C$ be a set of all important points such that $v_{c}=1,\leavevmode\nobreak\ \forall c\in\mathcal{C}_{+}$. Then other points are less important as $v_{c^{\prime}}\ll 1,\leavevmode\nobreak\ \forall c^{\prime}\in C\setminus\mathcal{C}_{+}$. The algorithm guarantees an optimal solution when agents can cover all the important points as follows.
Theorem 7.
If $N=|\mathcal{C}_{+}|$, Algorithm 2 converges to an optimal solution.
Proof.
Note that each agent is allocated to a point $c\in\mathcal{C}_{+}$ in an optimal solution. Now let us assume that the algorithm converges to a sub-optimal solution $(x,\mathcal{P})$. In this case, some agents including $i_{\min}$ do not have any points $c\in\mathcal{C}_{+}$ in their partitions, due to the neighborhood optimality. Then, there is at least one agent $i$ whose partition $\mathcal{P}_{i}$ includes more than two points in $\mathcal{C}_{+}$. This violates the condition (17), which must be satisfied in the convergent state $Z_{4}$. This is a contradiction. ∎
Note that $\mathcal{N}^{2}_{i}$ is introduced for the scalability of the algorithm and is not necessary for the results above. Theorem 5 and 7 hold with the naive definition of neighborhood $\mathcal{N}_{i}$. Note that when there are multiple candidates of $i_{\min}$, Algorithm 1 forms a forest that consists of disjoint trees rooted at those candidates, and the problem is split into independent subproblems for each communication tree. To ensure the theorems hold, we replace function Neighbor in Line 6 in Algorithm 2 with Algorithm 3 to synchronize the timing of updating the communication trees. The algorithm introduces the flag $s_{i}$ for synchronization. Agent $i$ is initialized with $s_{i}=False$ and updates it as $s_{i}=True$ when $i$ is picked up at Line 5 in Algorithm 2. With this flag, Algorithm 3 checks whether all the agents have been processed (Line 2). If yes, the algorithm resets the memory (Line 3) and tries to update the communication trees of all agents (Line 5). Otherwise, the algorithm tries to update each communication tree independently. To this end, the set of agents in the tree is identified (Line 7) and the flags of those agents are initialized (Line 8-9). $\textsc{TreeMember}(i)$ is a function to identify the member of the communication tree to which $i$ belongs. Then the procedure CommTree of $i$ is triggered to construct the communication trees (Line 10) and the algorithm returns the new neighborhood $\mathcal{N}^{2}_{i}$. This synchronization ensures that the entire communication forest is updated only when all the trees are processed once, thereby ensuring the progress of algorithm convergence.
Algorithm 3 Synchronized update of neighborhood
1:procedure $\mathcal{N}^{2}_{i}=$Neighbor($\mathcal{P},i$)
2:     if $s_{k}=True,\forall k\in\mathcal{N}$  then
3:         $s_{k}=False,\forall k\in\mathcal{N}$
4:     if $s_{k}=False,\forall k\in\mathcal{N}$ then
5:         $\hat{u}_{k}=u_{k},D_{k}=-1,\mathcal{\hat{N}}_{k}=\mathcal{N}_{k},\forall k\in\mathcal{N}$
6:     else
7:         $\mathcal{N}^{tree}_{i}=\textsc{TreeMember}(i)$
8:         $\hat{u}_{k}=u_{k},D_{k}=-1,\forall k\in\mathcal{N}^{tree}_{i}$
9:         $\mathcal{\hat{N}}_{k}=\mathcal{N}_{k}\cap\mathcal{N}^{tree}_{i},\forall k\in\mathcal{N}^{tree}_{i}$
10:     $D_{i}=\textsc{CommTree}(i)$
11:     $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$
12:     Return $\mathcal{N}^{2}_{i}$
3.2. Complexity Analysis
As mentioned previously, Algorithm 2 requires solving as a subroutine an NP-hard problem (to compute the neighborhood optimum $M_{n_{i}}$ in Line 10 of Algorithm 2). This computation is the bottleneck of the algorithm, and its computational complexity is $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{n_{i}})$: we need to compute for an agent the optimal way to allocate her neighbors within her partition $\mathcal{P}_{n_{i}}$. In the general case, the size of the neighborhood $n_{i}$ can be as high as $n-1$. This occurs when an agent is at the center of a circle and can be neighbor of all the agents located in that circle. This makes the algorithm non-scalable. Nevertheless, we introduce the neighborhood with the constant size ($n_{i}=2$), which makes the algorithm polynomial in $n$.
As for the communication complexity of the algorithm, we start by bounding the convergence rate, which is the number of iterations before convergence. Let $d_{\max}$ be the upper bound of the distance between any two points in the environment $C$, and $\Delta v$ be the resolution limit of the weight $v_{c}$. In the case of discrete settings, the potential function $\phi(x,\mathcal{P})$ consists of the elemental term $v_{c}\leavevmode\nobreak\ g(|x_{i}-c|)$ and then the improvement in the potential function for each iteration is lower bounded by $\epsilon=\Delta v\cdot g(d_{\max})$. (In the case of continuous settings, we can regard $\epsilon$ as the agents’ resolution limit of utility). Note that the convergence of the algorithm is guaranteed by Theorem 5. For each iteration, Algorithm 2 picks an agent one by one and checks if the potential function can be improved or not. Before convergence, at least one out of $n$ agents improves the potential function. Then the convergence rate $\alpha$ is bounded as $\alpha\leq n[\phi(x^{*},\mathcal{P}^{*})-\phi(x,\mathcal{P})]/\epsilon$, where $(x,\mathcal{P})$ and $(x^{*},\mathcal{P}^{*})$ are the initial allocation and the allocation after convergence, respectively.
For each iteration, the agents synchronize the start and the end timing of the step, share the global information, share their private information to compute a neighborhood optimum, and finally share the neighborhood optimum solution. Agents also need to share the timing of updating the communication trees in Algorithm 3. To this end, we define additional global variables $s_{True}$ and $s_{False}$ that $s_{True}=True$ if $s_{i}=True$ for all agents, $s_{False}=True$ if $s_{i}=False$ for all agents, otherwise they are $False$. The communications to share the global information require at most $\mathcal{O}(n^{2})$ messages. The message to share the global information $(u_{\min},i_{\max}^{+},x_{i_{\min}},s_{True},s_{False})$ has a fixed length regardless of the problem size. If we assume $1/g$ is a polynomial function, $1/\epsilon$ and $\alpha$ are also polynomial in the problem size, and then the communication complexity is also polynomial.
4. Experiments
In order to validate the practical efficiency and scalability of our approach, we ran simulations. First, we evaluate the efficiency with small graphs, then we evaluate the scalability with larger graphs.
In what follows, the nodes in an environment graph are classified into two groups, which are $c\in\mathcal{C}_{+}$ with $v_{c}=1$ and $c^{\prime}\in C\setminus\mathcal{C}_{+}$ with $v_{c^{\prime}}\ll 1$. Nodes in $\mathcal{C}_{+}$ and the initial position of agents are allocated uniformly at random in the environment graph. We run 32 simulations for each experiment. All the error bars in the figures show 95% confidence intervals. Note that the environment can be any dimensional space, even though all the graphs are projected into 2D figures. All the numerical results are summarized in Table 1. As for the implementation, we use Python 3.8.12, Red Hat Enterprise Linux Server release 7.9 and Intel Xeon CPU E5-2670 (2.60 GHz), 192 GB memory to run the experiments. The random number generator is initialized by numpy.random.seed at the beginning of the main code, with different seeds for each run of the simulation.
Comparison.
The neighborhood optimal approach proposed in Section 3.1, coined in the following as NBO, is compared to:
  • (VVP) the vanilla distributed covering algorithm based on Voronoi partitioning, as described in Section 2.5.
  • (SOTA) the algorithm of Sadeghi et al. Sadeghi et al. (2020). In a nutshell, for a given agent $i$, the algorithm first tries to perform individual moves to maximize the social welfare of her neighborhood $\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x)$. If no such move is improving, it considers coordinated moves with a single other agent $j$, in the sense that $i$ would move and $j$ would take the place of agent $i$. The algorithm first considers neighbors, and then (via neighborhood communication), may consider agents further away. However, these coordinated moves only involve two agents at most.
  • (CGR) the centralized greedy algorithm that allocates agents one by one starting from the environment with nobody.
We evaluate the performance of the algorithms above with different shapes of the environments (Figure 6). In addition of these shapes, we also use the small bridge setting and OR library dataset shown in Yun and Rus (2014). More details of the shapes are in the supplementary material.
Figure 6. Different shapes of the environment.
Figure 7. Efficiency ratio $G(x)/G(x^{*})$ where $x$ is a solution of each algorithm and $x^{*}$ is an optimal solution, with $N=5$.
4.1. Efficiency
First, we evaluate the efficiency of the proposed neighborhood optimum algorithm, by comparing it with an optimal solution and the benchmarks. The efficiency is measured with the ratio $G(x)/G(x^{*})$ where $x$ is a solution of the algorithm and $x^{*}$ is an optimal solution. Since finding an optimal solution is NP-hard, the simulation is conducted on a 1D chain where $|C|=20,|\mathcal{C}_{+}|=10$, and $N=5$.
In that case, the proposed algorithm outperforms both benchmarks and improves the efficiency by about 21.7% points compared to SOTA (Figure 7). The efficiency ratio of NBO is 90% which is much better than the theoretical approximation ratio of 2. We examine the detail with an illustrative example in the supplementary material.
Due to the NP-hardness, it is difficult to compare with optimal solutions in larger settings. Instead, we compute the efficiency ratio with the centralized greedy (CGR) as a basis. Table 1 summarizes all the results obtained. The quantitative comparison of the efficiency shows similar results as in Figure 7 that the proposed method outperforms the benchmarks in all cases. Note that NBO even outperforms CGR sometimes despite the following two disadvantages of NBO compared to CGR: 1. NBO uses only limited information while CGR uses full information, 2. NBO’s initial state is less favorable compared to CGR’s matroidal setting.
Shapes SOTA VVP NBO
1D Chains 0.686 ± 0.045 0.515 ± 0.052 0.903 ± 0.050
Stars 0.983 ± 0.008 0.962 ± 0.010 1.017 ± 0.006
Trees 0.952 ± 0.009 0.935 ± 0.009 0.980 ± 0.004
Simualed indoor 0.895 ± 0.007 0.841 ± 0.010 0.917 ± 0.013
Random ($w=1$) 0.932 ± 0.010 0.904 ± 0.008 0.963 ± 0.009
Random ($w=2$) 0.936 ± 0.008 0.901 ± 0.007 0.976 ± 0.009
Small bridge 0.975 ± 0.007 0.955 ± 0.012 1.000 ± 0.002
3D structure 0.960 ± 0.006 0.937 ± 0.007 0.982 ± 0.010
OR library 0.843 ± 0.126 0.963 ± 0.014 0.978 ± 0.006
Table 1. Comparison to VVP performance: Efficiency ratio $G(x)/G(x^{CGR})$ where $x$ is a solution of each algorithm and $x^{CGR}$ is a solution of CGR. Each cell shows mean ± standard deviation. Larger values mean better results.
4.2. Scalability
We evaluate the scalability of the proposed algorithm by changing the size of the space $|C|$ and the number of agents $n$. The results show that:
  • The runtime grows with a polynomial order of $|C|$.
  • The runtime decreases when $n$ increases because the size of partitions $|\mathcal{P}_{n_{i}}|$ also decreases.
This result closely aligns with the theoretical complexity $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{2})$ and demonstrates the applicability of the proposed method to real-world large-scale problems. Due to the space limit, we defer the details to the supplementary material.
5. Conclusion
To guarantee the performance of distributed coverage control, we extend the approach of Marden (2016) to general dimensional settings for discrete environments, by communicating additional information on the position of the minimum utility agent. The algorithm guarantees convergence to a neighborhood optimum solution, even in challenging non-convex settings and achieves the approximation ratio of 2, which is equivalent to the case of full information on all agents. This approximation ratio surpasses that of the state-of-the-art method presented in Sadeghi et al. (2020). While both the approach of Sadeghi et al. (2020) and ours are more communication demanding than simple best responses based on Voronoi partitioning, this remains manageable and prioritizes local interactions.
Furthermore, we show that the algorithm guarantees optimality in a special subclass of the coverage problem. Through the experiments, we illustrate that the proposed algorithm outperforms state-of-the-art benchmark algorithms. Additionally, runtime results confirm the scalability of the algorithm, consistent with theory.
Appendix A Proof of Theorem 5 (convergence)
For each iteration, Algorithm 2 will produce a new state $(x^{\prime},\mathcal{P}^{\prime})$ from an old state $(x^{\prime},\mathcal{P}^{\prime})$. Note that after step a,
$\begin{array}[]{l}M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (25)
Also, after step b,
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (28)
Lemma 0.
If $(x,\mathcal{P})\in Z_{1}$, then Algorithm 2 will produce a sequence of states that results in a new state $(x^{\prime},\mathcal{P}^{\prime})\in Z_{2}$.
Proof.
At first,
$\begin{array}[]{ll}V(x,\mathcal{P})&=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}M_{1}(x_{i};\mathcal{P}_{i})\\ &=M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ &=\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})-G(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ &=\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})-u_{i_{\max}^{+}}(x,\mathcal{P}).\end{array}$ (36)
In case of step a, if $i_{\min}(x,\mathcal{P})\in\mathcal{N}_{i}$,
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}}).\end{array}$ (44)
Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{n_{i}})\\ \geq\underset{\begin{subarray}{c}y_{1},\ldots,y_{k-2}\in\mathcal{P}_{n_{i}}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\max}G(y_{1},\ldots,y_{k-2};\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\setminus\{i_{\min}(x,\mathcal{P}),i_{\max}^{+}(x,\mathcal{P})\}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+\underset{\begin{subarray}{c}y\in\mathcal{P}_{i_{\max}^{+}}\end{subarray}}{\max}G(y,x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ =\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (58)
If $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$, since $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P})$,
$\begin{array}[]{ll}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})&=M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ &\geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-u_{\min}(x,\mathcal{P}).\end{array}$ (62)
Since $i_{\max}^{+}\in\mathcal{N}_{i}$, $\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})=\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})$ means a sum of utilities when $n_{i}+1$ agents share $\mathcal{N}_{i}$. Then,
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P}).\end{array}$ (65)
With (62), we have
$\begin{array}[]{l}\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x^{\prime},\mathcal{P}^{\prime})\geq\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (69)
In both cases, we have
$\begin{array}[]{lll}\phi(x^{\prime},\mathcal{P}^{\prime})&=&\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ &\geq&\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}\\ &&+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ &\geq&\phi(x,\mathcal{P}).\end{array}$ (76)
Note that the equality holds only when $V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})\leq 0$, which is $(x^{\prime},\mathcal{P}^{\prime})\in Z_{2}$. As long as the state stays in $Z_{1}$, $\phi$ increases.
In case of step b, $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$ and $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})>u_{\min}(x,\mathcal{P})$.
First, step b computes $(x^{\prime},\mathcal{P}^{\prime})$ as if an imaginary agent $l$ is added into $\mathcal{P}_{n_{i}}$, especially at $\mathcal{P}^{\prime}_{i_{+}}$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})=M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\end{array}$ (79)
Note that
$\begin{array}[]{ll}M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})\\ =M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ >u_{\min}(x,\mathcal{P}).\end{array}$ (85)
Meanwhile, since $i_{\max}^{+}\in\mathcal{N}_{i}$,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+V(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+M_{1}(x_{i_{\max}^{+}};\mathcal{P}_{i_{\max}^{+}})\\ \leq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}}).\end{array}$ (93)
Then, since $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$,
$\begin{array}[]{lll}\phi(x^{\prime},\mathcal{P}^{\prime})&=&\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ &=&\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x,\mathcal{P})]_{+}.\end{array}$ (98)
By the definition of $V$ and (85),
$\begin{array}[]{l}\geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})-u_{\min}(x,\mathcal{P}).\end{array}$ (101)
By (93),
$\begin{array}[]{l}\geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+V(x,\mathcal{P})-u_{\min}(x,\mathcal{P}).\end{array}$ (104)
Since $(x,\mathcal{P})\in Z_{1}$,
$\begin{array}[]{l}=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}=\phi(x,\mathcal{P}).\end{array}$ (107)
Note that if $V(x^{\prime},\mathcal{P}^{\prime})>M_{1}(x^{\prime}_{i_{+}},\mathcal{P}^{\prime}_{i_{+}})$, then $\phi(x^{\prime},\mathcal{P}^{\prime})>\phi(x,\mathcal{P})$. Also, from (85), $(x^{\prime},\mathcal{P}^{\prime})\in Z_{1}$.
[…]
Then, as long as $(x,\mathcal{P})\in Z_{1}$, $\phi(x,\mathcal{P})$ increases. Since the solution space is finite, $(x,\mathcal{P})$ reaches $Z_{2}$ in the end.
Lemma 0.
If $(x,\mathcal{P})\in Z_{2}\setminus Z_{3}$, then Algorithm 2 will produce a sequence of states that results in a new state $(x^{\prime},\mathcal{P}^{\prime})\in Z_{3}$.
Proof.
Since $(x,\mathcal{P})\in Z_{2}$, $V(x,\mathcal{P})\leq u_{\min}(x,\mathcal{P})$. In case of step a,
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}\\ =\phi(x,\mathcal{P}).\end{array}$ (120)
Note that the equality holds only when $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi$ increases as long as the local state in $\mathcal{N}_{i}$ does not satisfy the condition of $Z_{4}$.
In case of step b,
$\begin{array}[]{l}M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+M_{1}(x^{\prime}_{jnearest};\mathcal{P}^{\prime}_{jnearest})\\ \leq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime}).\end{array}$ (126)
Since $(x,\mathcal{P})\in Z_{2}\setminus Z_{3}$, $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})>u_{\min}(x,\mathcal{P})$. Then,
$\begin{array}[]{l}\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime})\\ \geq M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})\\ >M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+u_{\min}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+u_{\min}(x,\mathcal{P}).\end{array}$ (133)
Then, since $i_{\min}(x,\mathcal{P})\not\in\mathcal{N}_{i}$,
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x,\mathcal{P})]_{+}\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x,\mathcal{P})\\ >\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})+[V(x,\mathcal{P})-u_{\min}(x,\mathcal{P})]_{+}\\ =\phi(x,\mathcal{P}).\end{array}$ (146)
Then $\phi$ always increases as long as $(x,\mathcal{P})\in Z_{2}\setminus Z_{3}$. Note that $(x^{\prime},\mathcal{P}^{\prime})$ can be in $Z_{1}$. However, according to Lemma 8, the state will come back to $Z_{2}$ without cycle. Since $(Z_{2}\setminus Z_{3})\cup Z_{1}$ is finite, $(x,\mathcal{P})$ reaches $Z_{3}$ in the end. ∎
Lemma 0.
If $(x,\mathcal{P})\in Z_{3}$, then Algorithm 2 will produce a sequence of states that results in a new state $(x^{\prime},\mathcal{P}^{\prime})\in Z_{4}$.
Proof.
By definition,
$\begin{array}[]{l}M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y_{1},\ldots,y_{k}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{k};\mathcal{P}_{n_{i}})\\ \geq G(x;\mathcal{P}_{n_{i}}).\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \end{array}$ (153)
Since $(x,\mathcal{P})\in Z_{3}$, Algorithm 2 always runs step a. In this case, $V$ always decreases and $u_{\min}$ always increases. Then $(x,\mathcal{P}),(x^{\prime},\mathcal{P}^{\prime})\in Z_{2}$. Then,
$\begin{array}[]{l}\phi(x^{\prime},\mathcal{P}^{\prime})\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})+[V(x^{\prime},\mathcal{P}^{\prime})-u_{\min}(x^{\prime},\mathcal{P}^{\prime})]_{+}\\ =\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x^{\prime},\mathcal{P}^{\prime})\\ =M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})+\underset{\begin{subarray}{c}i\in\mathcal{N}\setminus\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ \geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})\\ =\phi(x,\mathcal{P}).\end{array}$ (164)
Note that the equality holds only when $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})$. This means that $\phi(x,\mathcal{P})$ increases as long as $(x,\mathcal{P})\in Z_{3}$, until $(x,\mathcal{P})$ reaches $Z_{4}$. ∎
Proof of Theorem 5 (Convergence).
It follows Lemma 8 to 10. ∎
Appendix B Proof of Theorem 5 (Approximation ratio)
Lemma 0.
If $(x;\mathcal{P})\in Z_{4}$, $M_{1}(x;\mathcal{P})\leq u_{\min}(x,\mathcal{P})$.
Proof.
Since $(x;\mathcal{P})\in Z_{4}\subseteq Z3$, $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Also, $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})=\underset{\begin{subarray}{c}i\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{i}(x,\mathcal{P})=G(x,\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall i\in\mathcal{N}$. Then, forall $i\in\mathcal{N}$,
$\begin{array}[]{l}M_{1}(x_{n_{i}},\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y,x_{n_{i}};\mathcal{P}_{n_{i}})-G(x_{n_{i}},\mathcal{P}_{n_{i}})\\ =\underset{\begin{subarray}{c}y\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y,x_{n_{i}};\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}x_{n_{i}}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(x_{n_{i}},\mathcal{P}_{n_{i}})\\ \leq\underset{\begin{subarray}{c}y_{1},\ldots,y_{n_{i}+1}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{n_{i}+1};\mathcal{P}_{n_{i}})-\underset{\begin{subarray}{c}y_{1},\ldots,y_{n_{i}}\in\mathcal{P}_{n_{i}}\end{subarray}}{\max}G(y_{1},\ldots,y_{n_{i}};\mathcal{P}_{n_{i}})\\ =M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\\ \leq u_{\min}(x,\mathcal{P}).\end{array}$ (176)
Then, $u_{\min}(x,\mathcal{P})\geq\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}M_{1}(x_{n_{i}};\mathcal{P}_{n_{i}})=M_{1}(x;\mathcal{P})$. ∎
Lemma 0.
$G(x;\mathcal{C})$ is submodular in $x$. For any allocation $x\in\mathcal{X}$, agent sets $S\subseteq T\subseteq\mathcal{N}$, and agent $i\in S$,
$\begin{array}[]{l}G(x_{S};\mathcal{C})-G(x_{S\setminus\{i\}};\mathcal{C})\geq G(x_{T};\mathcal{C})-G(x_{T\setminus\{i\}};\mathcal{C}).\end{array}$ (178)
Proof.
Because $vor_{i}(x_{S};\mathcal{C})\subseteq vor_{i}(x_{T};\mathcal{C})$. This is because of agents $T-S$. ∎
Lemma 0.
$M_{k}(x,\mathcal{C})$ is monotone decreasing in $x$.
$\begin{array}[]{l}M_{k}(x\setminus\{x_{i}\};\mathcal{C})\geq M_{k}(x;\mathcal{C}),\forall x_{i}\in\mathcal{C}.\end{array}$ (180)
Proof.
By Lemma 12,
$\begin{array}[]{l}G(y_{1},\ldots,y_{k},x\setminus\{x_{i}\};\mathcal{C})-G(x\setminus\{x_{i}\};\mathcal{C})\geq G(y_{1},\ldots,y_{k},x;\mathcal{C})-G(x;\mathcal{C}),\\ \forall y_{1},\ldots,y_{k}\in\mathcal{C}.\end{array}$ (183)
Proof of Theorem 9.
Let $(x,\mathcal{P})$ be a solution of Algorithm 2 and $(x^{*},\mathcal{P}^{*})$ be an optimal solution. Also, we denote $G(x;\mathcal{C})$ as $G(x)$. Then,
$\begin{array}[]{lll}G(x^{*})&\leq&G(x,x^{*})\\ &=&G(x)\\ &&+[G(x,x^{*}_{1})-G(x)]\\ &&+[G(x,x^{*}_{1},x^{*}_{2})-G(x,x^{*}_{1})]\\ &&+\ldots\\ &&+[G(x,x^{*})-G(x,x^{*}_{1},\ldots,x^{*}_{N-1})]\\ &\leq&G(x)\\ &&+M_{1}(x,\mathcal{P})\\ &&+M_{1}(x,x^{*}_{1},\mathcal{P})\\ &&+\ldots\\ &&+M_{1}(x,x^{*}_{1},\ldots,x^{*}_{N-1},\mathcal{P})\\ &\leq&G(x)+n*M_{1}(x,\mathcal{P})\\ &\leq&G(x)+n*u_{\min}(x,\mathcal{P})\\ &\leq&2G(x).\end{array}$ (198)
Appendix C Comparison of algorithms in 1 dimentional setting
We examine the different behavior of algorithms (VVP, SOTA, NBO and the optimal solution, OPT) with simple 1 dimentional setting (Figure 8). Initially, two agents are located at the left end. Agent 0 moves first in all cases. In SOTA, each agent is activated once. The agent first tries to improve the objective function by itself. Then, the agent also tries to communicate with other agents to improve the objective function together. In B, agent 0 tries to move first, but it can’t move because it is blocked by agent 1. Then, agent 0 is deactivated and does not move anymore by itself. Next, agent 1 moves right, stops at its best position, and tries to improve the objective function further by communicating with agent 0. However, the algorithm cannot find any collaborative moves to improve the situation and terminates. Meanwhile, in C, NBO finds an optimal solution because the two agents are neighbors. In this case, the coverage control reaches another optimal solution as in D.
As shown above, SOTA depends on the action order of agents. Then we ran the same simulation by switching the initial locations of the agents. In E, SOTA can move both agents, and they reach a better solution than B (but not optimal). Meanwhile, NBO can find an optimal solution in a stable manner as in F.
Figure 8. 1 dimentional setting for comparison of algorithms with $n=2,m=12$. All nodes have the same weight $v_{c}=1$. Agent 0 moves first in all results. A: The initial allocation. B …D: Outcomes of algorithms. $G_{*}$ is the value of the objective function. $G_{OPT}$ is the optimal value. E,F: Outcomes with switching the initial locations of the agents.
Appendix D The Details of Experiments
Shapes.
We ran our experiments on different shapes:
  • (1D) lines —While the original approach of Marden Marden and Wierman (2009) was sufficient on one-dimensional structures, we use it as a benchmark since optimal solutions can be computed when the number of agents remains reasonable.
  • (2D) stars and trees —We generated two representatives of these standard shapes (Figure 9).
  • (2D) simulated indoor environment —This is based on a possible application which could consist in building a mesh network in an indoor environment (Figure 10), where coverage could be needed after a catastrophic event (e.g. building inaccessible after a nuclear accident).
  • (2D) randomly generated structures —We start from a template structure with the parameter of corridor width $w\in\{1,2\}$ (see Figure 11). From such a template structure, a connected structure is then generated by randomly removing nodes.
  • (2D) a small bridge —This is an example of non-convex discrete environment shown in Yun and Rus (2014).
  • (3D) non-planar graphs —To illustrate how the approach can work in higher dimensions, we selected some non-planar graphs (Figure 12).
  • ($\geq$ 3D) OR library —This is a test dataset consisting of random non-planar graphs for $p$-median problem Beasley (1990).
Though we also try to evaluate the cooperative version of VVP, where each agent $i$ tries to maximize the social welfare of her neighborhood, we discontinue the experiment due to its lack of convergence in practice. (For example, it fails to converge in more than 60% of cases in 3D setting). We also omit the evaluation of the algorithm in Yun and Rus (2014), because it can be regarded as a variant of VVP by skipping the moves that could change $\mathcal{N}_{\mathcal{N}_{i}}$. In the cases of OR library, we have results only for 6 instances due to the large problem size.
Figure 9. Left: A star with extended branches. Right: A tree. Light gray nodes show targets ($v_{c}=1$), and dark gray nodes show agents.
Figure 10. Setting for indoor applications with $N=21$. The space consists of narrow corridors and small rooms. Top left: initial allocation, Bottom left: proposed method, Top right: benchmark Sadeghi et al. (2020).
Figure 11. Randomly generated structures. Left: the template structure with the parameter of corridor width $w$, with $L=3(w+1)-1$, and the length of the tail $S=2w+4$. Right: connected structure generated by randomly removing nodes.
Figure 12. Setting for 3D application with $N=18$. The space consists of narrow corridors. Top left: initial allocation, Bottom left: proposed method, Top right: benchmark Sadeghi et al. (2020).
We evaluate the scalability of the proposed algorithm by changing the size of the space $|C|$ and the number of agents $n$. Also, we set $|\mathcal{C}_{+}|=|C|$ (the top figure in Figure 13). The middle figure shows the runtime until the convergence with different $|C|$. The growing speed of the runtime closely matches the theoretical prediction of $\mathcal{O}(|\mathcal{P}_{n_{i}}|^{2})$, which is polynomial. The bottom figure shows the runtime with different $n$. Surprisingly, the runtime decreases when $n$ increases, because the size of the partitions $|\mathcal{P}_{n_{i}}|$ also decreases. Note that the current implementation uses a single CPU just for theoretical verification, and parallel computation for each agent can reduce the runtime further.
Figure 13. Up: The environment ($|C|=152$) to evaluate the scalability. Middle: Runtime of our algorithm (Section LABEL:sec:scale) until the convergence when chainging $|C|$ ($n=20$). Bottom: Runtime of our algorithm when chainging $n$ ($|C|=192$).
Appendix E Reproducibility Checklist
E.1. Algorithm
If the paper introduces a new algorithm, it must include a conceptual outline and/or pseudocode of the algorithm for the paper to be classified as CONVINCING or CREDIBLE. (CONVINCING)
E.2. Theoretical contribution
If the paper makes a theoretical contribution:
  1. (1)
    All assumptions and restrictions are stated clearly and formally (yes)
  2. (2)
    All novel claims are stated formally (e.g., in theorem statements) (yes)
  3. (3)
    Appropriate citations to theoretical tools used are given (yes)
  4. (4)
    Proof sketches or intuitions are given for complex and/or novel results (yes)
  5. (5)
    Proofs of all novel claims are included (yes)
For a paper to be classified as CREDIBLE or better, we expect that at least 1. and 2. can be answered affirmatively, for CONVINCING, all 5 should be answered with YES. (CONVINCING)
E.3. Data sets
If the paper relies on one or more data sets:
  1. (1)
    All novel datasets introduced in this paper are included in a data appendix (NA)
  2. (2)
    All novel datasets introduced in this paper will be made publicly available upon publication of the paper (NA)
  3. (3)
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations (NA)
  4. (4)
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available (NA)
  5. (5)
    All datasets that are not publicly available (especially proprietary datasets) are described in detail (NA)
Papers can be qualified as CREDIBLE if at least 3., 4,. and 5,. can be answered affirmatively, CONVINCING if all points can be answered with YES. (NA)
E.4. Experiments
If the paper includes computational experiments:
  1. (1)
    All code required for conducting experiments is included in a code appendix (yes)
  2. (2)
    All code required for conducting experiments will be made publicly available upon publication of the paper (yes)
  3. (3)
    Some code required for conducting experiments cannot be made available because of reasons reported in the paper or the appendix (NA)
  4. (4)
    This paper states the number and range of values tried per (hyper-)parameter during development of the paper, along with the criterion used for selecting the final parameter setting (yes)
  5. (5)
    This paper lists all final (hyper-)parameters used for each model/algorithm in the experiments reported in the paper (yes)
  6. (6)
    In the case of run-time critical experiments, the paper clearly describes the computing infrastructure in which they have been obtained (yes)
For CREDIBLE reproducibility, we expect that sufficient details about the experimental setup are provided, so that the experiments can be repeated provided algorithm and data availability (3., 5., 6.), for CONVINCING reproducibility, we also expect that not only the final results but also the experimental environment in which these results have been obtained is accessible (1., 2., 4.). (CONVINCING)